OpenStudy (anonymous):

4=-4(-2)+8 in standard form

OpenStudy (anonymous):

is this the complete question, it just looks unusual

OpenStudy (anonymous):

yeah it is an incomplete question

OpenStudy (anonymous):

I need 4=-4(-2)+8 to be written in standard form. Instead of written in point-slope form.

OpenStudy (anonymous):

Standard form: ax + by = c

OpenStudy (anonymous):

btw where are the variables???

OpenStudy (anonymous):

still incomplete

OpenStudy (anonymous):

The question above it was to figure out wat the y intercept was, I had 4=-4(-2)+ b then worked it out to get the answer to b. b=8. Does this help.

OpenStudy (anonymous):

slope-intercept form is y = mx + b where: m = slope b = y-intercept So you already found your answer. b = 8

OpenStudy (anonymous):

Yes but I need this equation to turn into standard form. Or is it already in standard form.

OpenStudy (anonymous):

What was the starting equation they gave you?

OpenStudy (anonymous):

Ok The starting equation was y+4=-4(x-2) then it asked me to rewrite it in slope-intercept form and then to figure out b. now it wants it in standard form.

OpenStudy (anonymous):

Before: 4 = -4(-2) + 8 To change into Standard form (ax + by = c) I need to have x and y.

OpenStudy (anonymous):

Original equation y+4=-4(x-2) x=-2y=4

OpenStudy (anonymous):

Okay so change y + 4 = -4(x - 2) into slope-intercept form: y + 4 = -4(x - 2) Distribute. Combine like terms and solve for y. y + 4 = -4x + 8 y = -4x + 4 So slope-intercept form would be: y = -4x + 4 But you want it in standard form: ax + by = c. y = -4x + 4 Just rearrange the order to: 4x + y = 4

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Yes, but now it's asking wat the xintercept is.

OpenStudy (anonymous):

To find x, make y = 0. To find y, make x = 0.

OpenStudy (anonymous):

So if you want to find x: 4x + y = 4 4x + (0) = 4 And solve.

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

Yes Ty.

OpenStudy (anonymous):

welcome. glad I could help (:

OpenStudy (anonymous):

Do you think you could help me with another one.